[clas_members] Status update and shifts

Eugene Pasyuk pasyuk at jlab.org
Thu Sep 20 09:27:50 EDT 2018

Dear Collaborators,

Here is the status update.
The accelerator is recovering from shutdown and there is a beam in the machine. Two problems which caused the shutdown, vacuum and obstruction are fixed. They are planning to lock up halls on swing shift today and start tuning beam lines for halls tomorrow. They hope to send physics quality beam this weekend.
On the Hall B side. All detectors brought back to operation after hurricane. The remaining problem is our cryo target.
PLC controlled boards came back from Saclay and they seem to be operational. However we could not read any sensors inside the cryostat. It was opened up and we found melted connectors inside the can. The connectors are being replaced now. The plans is to close it up today, pump down overnight and attempt to cool it down tomorrow. If it works we will reassemble beam line and lock up the hall. Hall will stay open until the target is operational. 

Shifts are canceled until 16:00 tomorrow, Friday September 21.

Stay tuned for further updates.


More information about the clas_members mailing list